OpenStudy (anonymous):

Aran moved a cone-shaped pile of sand that had a height of 6 ft and a radius of 3 ft. He used all of the sand to fill a cylindrical pit with a radius of 6 ft. How high did the sand reach in the pit? Use 3.14 to approximate pi and express your final answer in tenths. ft

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (anonymous):

@BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001 @BlossomCake @confluxepic @iGreen @SolomonZelman @TheSmartOne @shifuyanli @peterg2001

OpenStudy (igreen):

Don't tag so many times..that's against the Code of Conduct.

OpenStudy (igreen):

Find the volume of the sand. V = 1/3 * pi * r^2 * h V = 1/3 * 3.14 * 3^2 * 6 Simplify

OpenStudy (igreen):

Can you do that? @thebro55

OpenStudy (confluxepic):

Please don't tag me so many times in one post. Once is enough.

OpenStudy (anonymous):

56.52

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, now use the formula for the volume of a cylinder to find the height V = pi * r^2 * h 56.52 = 3.14 * 6^2 * h Simplify 3.14 * 6^2

OpenStudy (igreen):

@thebro55

OpenStudy (igreen):

Solve that equation for 'h' and you got your answer